OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function?

OpenStudy (anonymous):

choices are: -4, 12, -28, 60, … -4, -12, -28, -60, … -4, 12, -18, 54, … -4, 12, -18, -54, …

OpenStudy (anonymous):

@vangoo98 sorry to bother you but could you help me out on this one too please?

OpenStudy (anonymous):

I have no idea tbh o.e I haven't gotten to recursive functions yet. Sorry v.v What grade are you in btw?

OpenStudy (freckles):

try entering in 2 for n

OpenStudy (anonymous):

Its okay. do you know anyone that might be able to help me out on this one though? and I am a 10th grader but I am also taking 9th grade math because I fell behind in the school system @vangoo98

OpenStudy (freckles):

f(n) = -3•f(n -1) - 2•f(n - 2) f(2)= -3*f(2-1)-2*f(2-2) f(2)=-3*f(1)-2f(0) f(1) is given and so is f(0)

OpenStudy (anonymous):

Oh, I'm in 10th too, and I haven't gotten to this lol. And I gotcha And maybe @texaschic101

OpenStudy (freckles):

replace f(1) with 12 and replace f(0) with -4

OpenStudy (anonymous):

Thanks so much :) and @freckles I am confused on what you said :/

OpenStudy (freckles):

may i ask on which part

OpenStudy (anonymous):

The beginning.. I am really confused on how to do this

OpenStudy (freckles):

I replaced n with 2

OpenStudy (freckles):

because we didn't have f(2) yet but we have f(1) and f(0)

OpenStudy (anonymous):

so like what you had did: f(n) = -3•f(n -1) - 2•f(n - 2) f(2)= -3*f(2-1)-2*f(2-2) f(2)=-3*f(1)-2f(0) f(1) is given and so is f(0) so do we just plug in the f(1) and also the f(0)?

OpenStudy (freckles):

yes

OpenStudy (freckles):

replace f(1) with 12 and f(0) with -4

OpenStudy (freckles):

these values were given

OpenStudy (freckles):

f(2)=-3*12-2*(-4)

OpenStudy (anonymous):

so I take that ^ and solve it to get the answer?

OpenStudy (freckles):

your answers are given in the form f(0),f(1),f(2),f(3),... f(0) is the 1st term f(1) is the 2nd term f(2) is the 3rd term f(3) is the 4th term

OpenStudy (freckles):

so evaluating that will give you your third term

OpenStudy (anonymous):

so then would it be B then?

OpenStudy (freckles):

why do you think it is B?

OpenStudy (freckles):

Isn't your first two terms... -4,12, not -4,-12

OpenStudy (anonymous):

A?